# Rethinking Circuit Completeness in Language Models: AND, OR, and ADDER Gates

### v1

@@ 4.1 Separating AND, OR, and ADDER Gates
Figure 3: $\Delta KL$ in removing 1 and 2 edges of AND, OR, ADDER gates.
However, the intersection operation mentioned above raise concerns about the granularity alignment between $\mathcal{C}_{\text{Ns}}$ and $\mathcal{C}_{\text{Dn}}$. As illustrated in Figure 2, if the number of edges in $\mathcal{C}_{\text{Ns}}$ significantly exceeds that in $\mathcal{C}_{\text{Dn}}$, some edges identified as the AND gates could belong to the true type of the ADDER gates. This misalignment in granularity can also occur when $\mathcal{C}_{\text{Dn}}$ is considerably larger. Therefore, we propose two metrics (refer to Appendix E) to assess the degree of misalignment between $\mathcal{C}_{\text{Ns}}$ and $\mathcal{C}_{\text{Dn}}$ when performing intersection.
We report the misalignment of $\mathcal{C}_{\text{Ns}}$ and $\mathcal{C}_{\text{Dn}}$ at different scales in Appendix E. The results indicate that when the number of edges in the Dn circuit is approximately equal to that in the Ns circuit, both the misalignment score and its standard deviation reach an acceptable level. Therefore, throughout this paper, we assume that the optimal alignment occurs when Ns and Dn contain an equal number of edges and conduct experiments based on this assumption by scaling the number of edges identified by Ns and Dn strategies in a similar range.
4.2 Discovering Logically Complete Circuit
@@
4.3 Validation of Logically Complete Circuit
In this subsection, we focus on the faithfulness and completeness of the logically complete circuit (from our framework in Section 4.2, denoted by $\mathcal{C}_{\text{Ns+Dn}}$) and the circuit of existing work (since existing work generally adopts Ns as a basic intervention strategy, we denote it by $\mathcal{C}_{\text{Ns}}$). Similar to Section 4.1, we select GPT2-small as the computational graph, and ACDC, EAP, and EdgePruning as methods to represent greedy search, linear estimation, and differentiable mask, respectively. We examine the circuits obtained through Ns, Dn, and Ns+Dn. For instance, in ACDC, when intervening on each edge, we simultaneously compute the effect of substituting the clean activation with a corrupted one in the clean run, and the effect of substituting the corrupted activation with a clean one in the corrupted run. In EAP, we compute gradients under both clean and corrupted conditions. For EdgePruning, we replace Equation 1 with Equation 3 as the optimization objective. Detailed implementation can be found in Appendix D.1. These experiments are conducted on three mainstream tasks for circuit discovery, namely indirect object inference (IOI) (Wang et al., ), greater than (GT) (Hanna et al., 2023), and syntactic agreement (Yu et al., 2024). The details of these tasks are presented in Table 2.
@@ 4.3.1 Completeness
Following the definition of completeness in Section 2, we first compare the changes in KL divergence and accuracy for the corresponding tasks (IOI, GT, SA) when the circuit is removed from the computational graph. Specifically, we compare the differences between the original circuits (obtained through Ns) and the logically complete circuits (obtained through Ns+Dn) after removal, for three methods: ACDC, EAP, and EdgePruning. To account for the effects of sparsity, we constrain the number of edges in both circuits to remain consistent across six sparsity levels: 100, 200, 500, 1000, 2000, and 5000 edges.
@@
Figure 4 shows that, both in terms of KL divergence and accuracy, the performance of circuits removed through Ns+Dn is noticeably weaker compared to those removed through Ns. This corroborates Corollary 2, where we note that Ns, due to its inability to fully recover the OR gate, results in suboptimal completeness. Additionally, we observe that the gap between Ns and Ns+Dn is largest in both metrics in the EAP method (see the solid line in Figure 4), where Ns fails to recover any edges of the OR gate. Additionally, since ACDC and EdgePruning are generally able to identify one OR edge, the recovered OR edge exhibits some degree of randomness. In the case of ACDC, this randomness is influenced by the search order, while in the case of EdgePruning, it is influenced by the initial values of the mask.
@@ 4.3.2 Faithfulness
In Appendix F, we compare the circuits obtained using three strategies—Ns, Dn, and Ns+Dn—under the same sparsity constraints (specifically, we select edge counts of 100, 200, 500, 1000, 2000, and 5000) in terms of KL divergence and accuracy. The results show that, in terms of faithfulness, we have the relationship: $\mathcal{C}_{\text{Ns+Dn}}\approx\mathcal{C}_{\text{Ns}}>\mathcal{C}_{\text{Dn}}$. Figure 5 illustrates the average of the three methods on the IOI task to corroborate this conclusion. More results can be found in Figure 9 (a)-(c), which further supports the faithfulness requirements asserted in Corollary 2.
@@ 5.1 Graph Study
In Appendix G, we present the circuits of the AND, OR, and ADDER gates recovered by ACDC on the IOI task and map the functions of the components to those in the previous IOI circuit (Wang et al., ). We show parts of these circuits as demonstrations in Figure 6; each color represents one function in IOI circuit and blocks represent components at different locations, such as “a5.9” indicating the 9-th attention head in the 5-th layer, and “m8” referring to the MLP in the 8-th layer. The complete gate circuit can be found in Figure 10 of Appendix G. The results are interesting, revealing that the AND gates typically receive edges from different functions, suggesting that these functions must work together to support the receiver’s activation. In contrast, the OR gates almost exclusively receive edges from the same function, indicating that these edges are likely interchangeable due to their execution of the same function. The ADDER gates, on the other hand, tend to focus on combining two functions from different layers, with the activation generally considering the outputs of both shallow-layer and deep-layer functions.
@@ 5.2 Output Contribution
In Appendix H, we investigate the contribution of three types of logic gates to the output. The gate effect represents the contribution of the entire gate to the output and the edge effect represents the average contribution of each edge to the output. The results show that the contribution of the ADDER gates is significantly higher than that of the AND and OR gates. Furthermore, methods that focus on the edge effect, such as differentiable masks and linear estimation, lead to a higher average effect in the recovered circuit.
@@ 5.3 Proportion
In Appendix I, we present the number of AND, OR, and ADDER edges recovered by different methods. The results indicate that the proportion is closely related to the type of circuit discovery method used. For instance, greedy search selects all edges beyond the threshold, resulting in nearly equal numbers of AND, OR, and ADDER edges. In contrast, differentiable mask methods calculate the effect of each edge, which is disadvantageous for gates like AND and OR that contain multiple edges. As a result, the number of ADDER edges is significantly higher.
@@ 6 Conclusions
This paper systematically introduces three logic gates—AND, OR, and ADDER—to explain the essential requirements of circuit faithfulness and completeness. Furthermore, it provides an analysis of how existing circuit discovery methods perform with respect to these logic gates. Additionally, we propose an Ns&Dn-based method for separating the three logic gates, and for restoring a logically complete circuit. We empirically validate the differences in faithfulness and completeness between the logically complete circuit and existing circuits. Finally, we explore the relationships between the logic gates in terms of distribution, contribution, and functionality.
@@ 6.1 Limitations and Future Research
We acknowledge that the contribution of logically complete circuits to circuit research is largely concentrated in theoretical insights. That is to say, since there is no significant difference in faithfulness between the logically complete circuit ($\mathcal{C}_{Ns+Dn}$) and the existing circuit ($\mathcal{C}_{Ns}$), it is difficult to directly demonstrate its “superiority” over current work. Therefore, we would like to propose some future research directions below to reveal the potential contributions of logical gates in further model control.
A logically complete circuit provides a more granular and logically coherent perspective on the interpretability of a circuit. With a complete understanding of the logical relationships between edges, the circuit becomes more useful for offering insights into model control. For instance, when combined with sparse autoencoders (SAE), a logically complete circuit for different tasks can reflect whether the skills associated with these tasks can be combined. That is, if the circuit for task A requires the presence of $i\to j$, and the circuit for task B requires the removal of $i\to j$, knowing that $i\to j$ exists as an OR edge in task A resolves the conflict between the two circuits. Thus, a logically complete circuit offers a novel approach for verifying the potential combination of tasks through boolean satisfiability, which is treated as our future study.
@@ 6.2 Societal and Ethical Impact
Our work aims to facilitate the process of understanding and explaining the logical connections in language models, which is crucial for their continued safe development and deployment. We do not foresee logically complete circuit and logical gates being used towards adverse societal or ethical ends.
Appendix A How Does Circuit Logic Model Intervention?
A.1 AND Gate
@@ A.3 ADDER Gate
Appendix B How Does Circuit Logic Affect Faithfulness, Completeness, and Sparsity?
B.1 Faithfulness
@@ B.2 Completeness
Appendix C Validation of Logical Gates
C.1 Toy Model
Motivated by [Conmy et al., 2023], to study a toy transformer model with an AND, OR, and ADDER gates, we take a 1-Layer transformer model with two heads per layer, ReLU-based activations, and model dimension 1. Specifically, as shown in Figure 7, Let $A_{1}$ and $A_{2}$ be two attention heads with respective biases $\text{bias}_{1}$ and $\text{bias}_{2}$, both set to 1. The activation function $m$ is based on the ReLU nonlinearity. To ensure that the output of each attention head corresponds directly to its bias, we use a zero tensor as the input. For corrupted activations, we employ zero ablation—i.e., we directly remove the activations along the corresponding edges.
@@ C.1 Toy Model
Under these configurations, we evaluate the performance of existing methods on the toy model, as summarized in Table 1. For example, under the default Ns, ACDC [Conmy et al., 2023] (representing greedy search), EAP [Syed et al., 2024] (representing linear estimation), and EdgePruning [Bhaskar et al., 2024] (representing differentiable mask) all successfully identify both $A_{1}$ and $A_{2}$ in the AND and ADDER gates. However, in the OR gate, ACDC and EdgePruning identify only one of $A_{1}$ or $A_{2}$—the specific result depends on the search order in ACDC and the initialization of the mask in EdgePruning—while EAP fails to identify any high-effect edge.
@@
According to Corollary 1, Ns can at least guarantee the full recovery of AND and ADDER gates. Greedy search, by retaining previous steps with removed results, can identify one OR edge (a detailed analysis is provided in Appendix A). Differentiable mask, optimizing for faithfulness (as per Corollary 2), ensures that at least one OR gate is included (otherwise, optimal faithfulness cannot be achieved), while additional OR edges conflict with the sparsity constraint and are therefore removed. However, linear estimation, when computing the effect of each edge, keeps all other edges in their non-removed state, which results in a failure to detect the OR gate. For example, when computing the effect of the edge $A\to C$ in the OR gate $A\to C\leftarrow B$, the edge $C\leftarrow B$ is also in the clean activation state, leading to a very small effect for $A\to C$. Similarly, when computing the effect of $C\leftarrow B$, the edge $A\to C$ remains in the clean activation state. In summary, linear estimation is unable to detect any edges of the OR gate.
@@
Appendix D Experiment Details
D.1 Baselines
In this work, we select ACDC [Conmy et al., 2023] to represent greedy search methods, EAP [Syed et al., 2024] to represent linear estimation methods, and EdgePruning [Bhaskar et al., 2024] to represent differentiable mask methods. In the following sections, we provide a detailed exposition of the original design of each method under the Ns. strategy, the corresponding formulation under the Dn. strategy, and the final approach that integrates both—Ns.+Dn.—for recovering logically complete gates.
D.1.1 Greedy Search Example: ACDC

### v2

@@ 3.3 Granularity Alignment between $\mathcal{C}_{\text{Ns}}$ and $\mathcal{C}_{\text{Dn}}$
The intersection operation mentioned above raise concerns about the granularity alignment between $\mathcal{C}_{\text{Ns}}$ and $\mathcal{C}_{\text{Dn}}$. As illustrated in Figure 3, if the number of edges in $\mathcal{C}_{\text{Ns}}$ significantly exceeds that in $\mathcal{C}_{\text{Dn}}$, some edges identified as the AND gates could belong to the true type of the ADDER gates. This misalignment in granularity can also occur when $\mathcal{C}_{\text{Dn}}$ is considerably larger. Therefore, we propose two metrics (refer to Appendix F) to assess the degree of misalignment between $\mathcal{C}_{\text{Ns}}$ and $\mathcal{C}_{\text{Dn}}$ when performing intersection.
We report the misalignment of $\mathcal{C}_{\text{Ns}}$ and $\mathcal{C}_{\text{Dn}}$ at different scales in Appendix F. The results indicate that when the number of edges in the Dn circuit is approximately equal to that in the Ns circuit, both the misalignment score and its standard deviation reach an acceptable level. Therefore, throughout this paper, we assume that the optimal alignment occurs when Ns and Dn contain an equal number of edges and conduct experiments based on this assumption by scaling the number of edges identified by Ns and Dn strategies in a similar range.
4 Discovering Logically Sound Circuit
4.1 Optimization for Logically Sound Circuit
@@
4.2 Validation of Logically Sound Circuit
In this subsection, we focus on the faithfulness and completeness of the logically sound circuit (from our framework in Section 4.1, denoted by $\mathcal{C}_{\text{Ns+Dn}}$) and the circuit of existing work (since existing work generally adopts Ns as a basic intervention strategy, we denote it by $\mathcal{C}_{\text{Ns}}$). Similar to Section 3.1, we select GPT2-small as the computational graph, and ACDC, EAP, and EdgePruning as methods to represent greedy search, linear estimation, and differentiable mask, respectively. We examine the circuits obtained through Ns, Dn, and Ns+Dn. For instance, in ACDC, when intervening on each edge, we simultaneously compute the effect of substituting the clean activation with a corrupted one in the clean run, and the effect of substituting the corrupted activation with a clean one in the corrupted run. In EAP, we compute gradients under both clean and corrupted conditions. For EdgePruning, we replace Equation 1 with Equation 3 as the optimization objective. Detailed implementation can be found in Appendix E.3. These experiments are conducted on three mainstream tasks for circuit discovery, namely indirect object inference (IOI) (wanginterpretability, ), greater than (GT) (hanna2023does, ), and syntactic agreement (yu2024functional, ). The details of these tasks are presented in Table 2.
@@ 4.2.1 Completeness
Following the definition of completeness in Section 2, we first compare the changes in KL divergence and accuracy for the corresponding tasks (IOI, GT, SA) when the circuit is removed from the computational graph. Specifically, we compare the differences between the original circuits (obtained through Ns) and the logically sound circuits (obtained through Ns+Dn) after removal, for three methods: ACDC, EAP, and EdgePruning. To account for the effects of sparsity, we constrain the number of edges in both circuits to remain consistent across six sparsity levels: 100, 200, 500, 1000, 2000, and 5000 edges.
@@
Figure 4 shows that, both in terms of KL divergence and accuracy, the performance of circuits removed through Ns+Dn is noticeably weaker compared to those removed through Ns. This corroborates Corollary 1, where we note that Ns, due to its inability to fully recover the OR gate, results in suboptimal completeness. Additionally, we observe that the gap between Ns and Ns+Dn is largest in both metrics in the EAP method (see the solid line in Figure 4), where Ns fails to recover any edges of the OR gate. Additionally, since ACDC and EdgePruning are generally able to identify one OR edge, the recovered OR edge exhibits some degree of randomness. In the case of ACDC, this randomness is influenced by the search order, while in the case of EdgePruning, it is influenced by the initial values of the mask. Moreover, we show the detailed results including Dn in Appendix G.
@@ 4.2.2 Faithfulness
In Appendix H, we compare the circuits obtained using three strategies—Ns, Dn, and Ns+Dn—under the same sparsity constraints (specifically, we select edge counts of 100, 200, 500, 1000, 2000, and 5000) in terms of KL divergence and accuracy. The results show that, in terms of faithfulness, we have the relationship: $\mathcal{C}_{\text{Ns+Dn}}\approx\mathcal{C}_{\text{Ns}}>\mathcal{C}_{\text{Dn}}$. Figure 5 illustrates the average of the three methods on the IOI task to corroborate this conclusion. More results can be found in Figure 9 (a)-(c), which further supports the faithfulness requirements asserted in Corollary 1.
@@ 5.1 Graph Study
In Appendix I, we present the circuits of the AND, OR, and ADDER gates recovered by ACDC on the IOI task and map the functions of the components to those in the previous IOI circuit (wanginterpretability, ). We show parts of these circuits as demonstrations in Figure 6; each color represents one function in IOI circuit and blocks represent components at different locations, such as “a5.9” indicating the 9-th attention head in the 5-th layer, and “m8” referring to the MLP in the 8-th layer. The complete gate circuit can be found in Figure 10 of Appendix I. The results are interesting, revealing that the AND gates typically receive edges from different functions, suggesting that these functions must work together to support the receiver’s activation. In contrast, the OR gates almost exclusively receive edges from the same function, indicating that these edges are likely interchangeable due to their execution of the same function. The ADDER gates, on the other hand, tend to focus on combining two functions from different layers, with the activation generally considering the outputs of both shallow-layer and deep-layer functions.
@@ 5.2 Output Contribution
In Appendix J, we investigate the contribution of three types of logic gates to the output. The gate effect represents the contribution of the entire gate to the output and the edge effect represents the average contribution of each edge to the output. The results show that the contribution of the ADDER gates is significantly higher than that of the AND and OR gates. Furthermore, methods that focus on the edge effect, such as differentiable masks and linear estimation, lead to a higher average effect in the recovered circuit.
@@ 5.3 Proportion
In Appendix K, we present the number of AND, OR, and ADDER edges recovered by different methods. The results indicate that the proportion is closely related to the type of circuit discovery method used. For instance, greedy search selects all edges beyond the threshold, resulting in nearly equal numbers of AND, OR, and ADDER edges. In contrast, differentiable mask methods calculate the effect of each edge, which is disadvantageous for gates like AND and OR that contain multiple edges. As a result, the number of ADDER edges is significantly higher.
@@ 6 Conclusions
This paper systematically introduces three logic gates—AND, OR, and ADDER—to explain the essential requirements of circuit faithfulness and completeness. Furthermore, it provides an analysis of how existing circuit discovery methods perform with respect to these logic gates. Additionally, we propose an Ns&Dn-based method for separating the three logic gates, and for restoring a logically sound circuit. We empirically validate the differences in faithfulness and completeness between the logically sound circuit and existing circuits. Finally, we explore the relationships between the logic gates in terms of distribution, contribution, and functionality.
@@ 6.1 Limitations and Future Research
The three logical gates under discussion in this paper are, in principle, derived from the intersection of $\mathcal{C}_{Ns}$ and $\mathcal{C}_{Dn}$. Nevertheless, it must be acknowledged that other types of logical gates may also exist—for instance, the XOR gate—whose underlying logic cannot be captured solely through the simple intersection of $\mathcal{C}_{Ns}$ and $\mathcal{C}_{Dn}$. That said, we contend that the three gates obtained from this intersection—AND, OR, and ADDER—are already sufficient to cover all edges of the circuits constructed from $\mathcal{C}_{Ns}$ and $\mathcal{C}_{Dn}$. For the purposes of the present scope of research, this coverage is adequate.
Moreover, a logically sound circuit provides a more granular and logically coherent perspective on the interpretability of a circuit. With a complete understanding of the logical relationships between edges, the circuit becomes more useful for offering insights into model control. For instance, a logically sound circuit for different tasks can reflect whether the skills associated with these tasks can be combined. That is, if the circuit for task A requires the presence of $i\to j$, and the circuit for task B requires the removal of $i\to j$, knowing that $i\to j$ exists as an OR edge in task A resolves the conflict between the two circuits. Thus, a logically sound circuit offers a novel approach for verifying the potential combination of tasks through boolean satisfiability, which is treated as our future study. We have demonstrated the potential contributions of completeness research with a toy task of model unlearning, as shown in Appendix L.
@@ 6.2 Societal and Ethical Impact
Our work aims to facilitate the process of understanding and explaining the logical connections in language models, which is crucial for their continued safe development and deployment. We do not foresee logically sound circuit and logical gates being used towards adverse societal or ethical ends.
Appendix A The Preference between Two Types of Completeness Metrics
The earlier completeness evaluation method (proposed by [16]) in fact measures incompleteness, defined as $D(C\setminus K\,\|\,G\setminus K)$, where a lower score indicates better performance. In contrast, the current method [9] measures completeness, defined as $D(G\setminus C\,\|\,G)$, where a higher score is preferable. Importantly, these two metrics are not directly interchangeable — that is, $1-D(C\setminus K\,\|\,G\setminus K)\neq D(G\setminus C\,\|\,G)$.
Nevertheless, a comparison can still be made in terms of the standard deviation. It has been theoretically demonstrated that the earlier completeness method suffers from high variance due to insufficient sampling, leading to instability in the final results. To illustrate this, we provide a simple comparison between the previous and current evaluation methods, as shown in the Table 4:
Table 4: The standard deviation between two metrics of completeness.
Method 5 sampled K 10 sampled K 20 sampled K 30 sampled K New metric
ACDC 1.1±1.6 1.8±1.2 1.4±1.1 1.2±0.6 1.3±0.13
EAP 1.7±1.5 1.3±1.6 1.6±1.2 1.3±0.7 1.4±0.11
Edge-P 1.5±1.8 1.4±1.9 1.4±1.6 1.6±0.6 1.2±0.08
Ours 0.4±0.3 0.6±0.4 0.3±0.3 0.3±0.2 1.8±0.03
We conducted five different runs of circuit discovery using random seeds, with the random subset $K$ chosen from subcircuits of sizes ranging from 2 to 5 nodes. For the previous evaluation metric, we performed 5, 10, 20, and 30 sampling iterations.
The experimental results clearly support two conclusions:
1. For algorithms that inherently lack completeness (such as ACDC, EAP, and Edge-Pruning), as well as for our own method, which ensures completeness, the previous evaluation metric exhibits large standard deviations across all trials. This indicates that the sampling process for the previous metric is insufficient, rendering the results highly unreliable.
2. The new evaluation metric significantly reduces variance for all types of circuits, as it does not suffer from sampling issues. The variance observed in the results arises from minor differences in the circuits generated by different random seeds.
Appendix B How Does Circuit Logic Model Intervention?
B.1 AND Gate
@@ B.3 ADDER Gate
Appendix C How Does Circuit Logic Affect Faithfulness, Completeness, and Sparsity?
C.1 Faithfulness
@@ C.2 Completeness
Appendix D Validation of Logical Gates
D.1 Toy Model
Motivated by [2], to study a toy transformer model with an AND, OR, and ADDER gates, we take a 1-Layer transformer model with two heads per layer, ReLU-based activations, and model dimension 1. Specifically, as shown in Figure 7, Let $A_{1}$ and $A_{2}$ be two attention heads with respective biases $\text{bias}_{1}$ and $\text{bias}_{2}$, both set to 1. The activation function $m$ is based on the ReLU nonlinearity. To ensure that the output of each attention head corresponds directly to its bias, we use a zero tensor as the input. For corrupted activations, we employ zero ablation—i.e., we directly remove the activations along the corresponding edges.
@@ D.1 Toy Model
Under these configurations, we evaluate the performance of existing methods on the toy model, as summarized in Table 1. For example, under the default Ns, ACDC [2] (representing greedy search), EAP [8] (representing linear estimation), and EdgePruning [10] (representing differentiable mask) all successfully identify both $A_{1}$ and $A_{2}$ in the AND and ADDER gates. However, in the OR gate, ACDC and EdgePruning identify only one of $A_{1}$ or $A_{2}$—the specific result depends on the search order in ACDC and the initialization of the mask in EdgePruning—while EAP fails to identify any high-effect edge.
@@
Ns can at least guarantee the full recovery of AND and ADDER gates. Greedy search, by retaining previous steps with removed results, can identify one OR edge (a detailed analysis is provided in Appendix B). Differentiable mask, optimizing for faithfulness (as per Corollary 1), ensures that at least one OR gate is included (otherwise, optimal faithfulness cannot be achieved), while additional OR edges conflict with the sparsity constraint and are therefore removed. However, linear estimation, when computing the effect of each edge, keeps all other edges in their non-removed state, which results in a failure to detect the OR gate. For example, when computing the effect of the edge $A\to C$ in the OR gate $A\to C\leftarrow B$, the edge $C\leftarrow B$ is also in the clean activation state, leading to a very small effect for $A\to C$. Similarly, when computing the effect of $C\leftarrow B$, the edge $A\to C$ remains in the clean activation state. In summary, linear estimation is unable to detect any edges of the OR gate.
@@
Appendix E Experiment Details
E.1 Why Does Search Strategy by repeating interventions on combinations is NP Problem?
Due to the differences in circuit algorithms, this issue cannot be fully proven mathematically. However, we will provide some additional explanation: we treat each ablation as $O(1)$. Assuming the language model contains $N$ nodes, if we need to use all possible intervention combinations to obtain a complete circuit, the time complexity would be $O(2^{N})$, which far exceeds polynomial time. However, different strategies can optimize the solution verification process to varying degrees. For example, EAP only requires $O(N)$ time to verify whether the circuit is complete, while Edge-pruning involves gradient descent, making its time complexity difficult to estimate. Additionally, greedy search can only perform a greedy verification in $O(N)$ time rather than a thorough verification. Therefore, we ultimately classify this problem as NP problem.
However, our proposed method, which employs the algorithm $N_{s}+D_{n}$, does not introduce additional computational complexity to the existing circuit discovery algorithm. Conceptually, it is equivalent to reapplying the $N_{s}$ algorithm once more under the strategy of $D_{n}$. As a result, the overall time complexity increases by at most a factor of two relative to the baseline algorithm, without imposing any additional nonlinear burden.
E.2 Detailed Results of Figure 3
Table 5: The changes in KL divergence between the original output and the new output when one or two edges are ablated.
AND OR ADDER
edge1(state) edge2(state) ($\Delta$ KL)(state) edge1(state) edge2(state) ($\Delta$ KL)(state) edge1(state) edge2(state) ($\Delta$ KL)(state)
remain (1) remain(1) 0.50 (1) remain (1) remain(1) 0.28 (1) remain (1) remain(1) 0.65 (2)
remain (1) ablate(0) 0.02 (0) remain (1) ablate(0) 0.27 (1) remain (1) ablate(0) 0.32 (1)
ablate (0) remain(1) 0.04 (0) ablate (0) remain(1) 0.27 (1) ablate (0) remain(1) 0.36 (1)
ablate (0) ablate(0) 0.00 (0) ablate (0) ablate(0) 0.00 (0) ablate (0) ablate(0) 0.00 (0)
All three subcircuits exhibit a fundamental pattern: the effect on the output is maximized when both edges are retained, and it drops to zero when both edges are ablated (as there is no clean activation left to support the node in this case). However, ablation of a single edge yields divergent outcomes:
In AND gate, ablating either edge nearly eliminates the effect on the output. In OR gate, ablating either edge has almost no impact on the output. In ADDER gate, ablating either edge leads to a noticeable degradation in the output.
E.3 Baselines
In this work, we select ACDC [2] to represent greedy search methods, EAP [8] to represent linear estimation methods, and EdgePruning [10] to represent differentiable mask methods. In the following sections, we provide a detailed exposition of the original design of each method under the Ns. strategy, the corresponding formulation under the Dn. strategy, and the final approach that integrates both—Ns.+Dn.—for recovering logically complete gates.
E.3.1 Greedy Search Example: ACDC